OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = 7x2 + 21x – 1 Question 3 answers (-2, -15) (2, 69) (1, 27) (-3/2, -67/4)

OpenStudy (experimentx):

for vertex use formula x = -b/2a = -21/14 => -3/2 your vertex would be the last one